OpenStudy (meilendurcer):

A salesperson receives a salary of $200 a week and makes a commission of 4% on all sales. Find the salesperson’s total earnings in a week with $3,000 in sales.

OpenStudy (anonymous):

200+.04(3000)=?

OpenStudy (lilg132):

3000 / 100 = 30 (this is 1%) multiply this by 4 to get 4% 30 * 4 = 120 add this to his salary 200 + 120 = 320 dont forget to click best response !